A researcher determines that students study an average of 8020 (MSD) minutes per week. Assuming these data are normally distributed, what is the z score for students studying 60 minutes per week? Show your work on the paper provided.
Answer #1

Solution :

Given that ,

mean = = 80

standard deviation = = 20

x = 60

z = (x - ) / = (60 - 80) / 20 = -20 / 20 = -1

z score = -1